Find P(-1.2 <2<-0.2)


1
Expert's answer
2022-05-22T23:55:15-0400

The probability of P(–a < Z < –b) is illustrated below:



First separate the terms as the difference between z-scores:

"P(-a<Z<-b)=P(Z<-b)-P(Z<-a)"

Then express these as their respective probabilities under the standard normal distribution curve:

"P(Z<-b)-P(Z<-a)=\Phi(-b)-\Phi(-a)"

"=(1-\Phi(b))-(1-\Phi(a))"

"=1-\Phi(b)-1+\Phi(a)"

"=\Phi(a)-\Phi(b)"

The above calculations can also be seen clearly in the diagram below:



So,

"P(-1.2<Z<-0.2)=\Phi(1.2)-\Phi(0.2)=0.8849-0.5793=0.3056" (using Z-score table)